Two charges are some distance apart and experience an electric force of 48 Newtons. What is the magnitude of the electric force if the charges do not change, but the distance between them is increased by a factor of 4.4? Answer in Newtons.
Answer #1

ス 니오 しス 2 Fr 4 8 G-4) 19-36
Answer #2

SOLUTION :


F is inversely proportional to  squared distance.

=> F = k / d^2 

So,


F1 = k  / (d1)^2 = 48 N

=> k = 48 (d1)^2 

F2 = k / (d2)^2

=> F2 = [48 *(d1)^2]  / (d2)^2   = 48 / (d2 / d1)^2  = 48 / 4.4^2  =  2.48 N


Hence, force between charges will be 2.48 N if distance is increased 4.4 times (ANSWER).

answered by: Tulsiram Garg